# Performance Gap in Entity Knowledge Extraction Across Modalities in Vision Language Models

### v2

@@ 1 Introduction
Over the past several years, vision-language models (VLMs) have made remarkable progress in linking visual inputs with natural language descriptions, enabling them to perform tasks such as image recognition Radford et al. (2021a); Mirza et al. (2024), video understanding Lin et al. (2024), and Visual Question Answering (VQA) Liu et al. (2023b, a). Much of the existing research evaluating these models focuses on their capacity to understand and describe information presented in the input image: counting objects, assessing complex visual arrangements, and understanding interactions between humans and objects Dang et al. (2024); Basu et al. (2024). While this has demonstrated that VLMs can integrate visual and textual information effectively, it does not fully address a critical question:
@@
Our empirical findings reveal a large accuracy drop when relying on visual representations rather than textual ones. Further probing suggests that this stems from the model investing too many layers in identifying the subject visually, leaving fewer layers for the second hop of factual reasoning (Biran et al., 2024). To investigate these dynamics more concretely, we utilize techniques of mechanistic interpretability such as activation patching (Vig et al., 2020) and attention knockout (Geva et al., 2023). They allow us to intervene in the model’s processing mid-inference, altering how and when image information influences the final answer. By tracking the flow of information through the network, we show that entity recognition and knowledge extraction are deeply entangled processes, with identification dominating the representational space longer than what one would expect. Consequently, the model’s internal knowledge is leveraged less efficiently.
@@ 2 Problem Setup
VLMs are architectures that integrate visual and textual modalities to perform tasks such as image captioning, visual question answering and more. Our VLM analysis in this work is mainly focused on the common architecture of LLaVA Liu et al. (2023b), and thus, we will follow the same notation.
@@
Every token in this input can be written as $\mathbf{h}_{i,v}^{0}$ or $\mathbf{h}_{i,t}^{0}$ where the $v$ and $t$ indicate a visual or textual token, the $i$ indicates the $i^{\text{th}}$ visual or textual tokens, and the 00 indicates input to layer 00. As the language model processes the entire input at each layer $l$, it builds a representation $\left(\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell},\mathbf{h}_{1,t}^{\ell}\dots\mathbf{h}_{m,t}^{\ell},\mathbf{h}_{1,g}^{\ell}\dots\mathbf{h}_{k,g}^{\ell}\right)$, where $\mathbf{h}_{1,g}^{\ell}...\mathbf{h}_{k,g}^{\ell}$ are the generated tokens hidden representations, $\left|\mathbf{H}_{\mathbf{v}}\right|=n$, $\left|\mathbf{H}_{\mathbf{t}}\right|=m$ and $\ell$ is the layer. We refer to the tokens once inside the LLM as hidden states.
@@
Through activation patching (Vig et al., 2020) and other mechanistic interpretability methods, we trace how and where entity recognition occurs, and how it impacts the model’s ability to retrieve the correct facts. Section 5 describes how we employ activation patching to identify the layer at which the model extracts the entity identity, establishing an upper bound on identification layers. Building on these findings, Section 6 applies forward activation patching to investigate whether the entity identity could be inferred at earlier layers.
We include additional experiments using attention knockouts (Geva et al., 2023) in appendix B, which further validate our conclusions. Throughout the paper, we use POPVQA to systematically evaluate VLM performance on visually driven factual retrieval and demonstrate how visual and textual components interact within a unified transformer architecture.
@@ 3 The POPVQA Dataset
In order to isolate the model’s ability to extract knowledge about a given entity from its ability to identify it in the first place, we want to restrict our questions to entities that the model recognizes. Other datasets that have knowledge seeking goals such as infoSeek (Chen et al., 2023) or VeQuAE (Lerner et al., 2022) do not include the entity in the image but rather only the answer to the question, thus, making it difficult to do this separation. Therefore, we built the POPVQA dataset to facilitate this need. It is inspired by PopQA Mallen et al. (2023) that created a QA dataset for various entities using Wikipedia.
The first part of POPVQA comprises of 15,395 entity-image pairs from Wikidata (Vrandečić and Krötzsch, 2014) along with aliases and popularity score (number of page views). The vast majority of entities are celebrities (63.6%). To allow ablation of other entity types, we added landmarks (18%), such as famous structures or UNESCO sites, famous paintings (14.6%), and brand logos (3.8%). All images are resized to 336 × 336 which is a standard input resolution, by adding black borders to create a square, then resizing. This was done in order to maintain the original aspect ratio, without losing any details. The popularity score was used only in the creation of the dataset such that we select only entities with high popularity, which increases the chance that the models will recognize them Mallen et al. (2023). We do not use this value directly in the experiments.
[…]
@@ 4 Motivating Experiment
We show results for the following models; LLaVA-1.5-7B (Liu et al., 2023b) as the canonic base VLM; we refer to it as LLaVA7B. LLaVA-MORE-8B (Cocchi et al., 2025) which uses a larger Llama 3.1 as its LLM and we try both variants using CLIP (Radford et al., 2021b) and SigLIP (Zhai et al., 2023) as backbones allowing for controlled comparison between visual backbones; we refer to these as LM-CLIP and LM-SigLIP. We use LLaVA-1.6-34B to assess the effect of model scale, and refer to it as LLaVA34B. We also compare to Qwen2-VL (Wang et al., 2024) that has a different architecture and training scheme.
@@ 7 Related Work
While VLMs demonstrate impressive capabilities in tasks such as visual reasoning, object recognition, and multimodal understanding Basu et al. (2024), their internal mechanisms and interpretability remain active areas of research. Below, we review the related works categorized into general VLM interpretability, cross-modal information flow, attention mechanisms, and mechanistic approaches to visual token processing.
@@
Dang et al. (2024) survey VLMs proposing a framework categorizing interpretability across data, model, and training perspectives. They highlight the alignment and decomposition of modalities as key challenges, emphasizing the need for explainability to ensure trustworthiness in multimodal AI applications. In this context, techniques for token-level, neuron-level, and embedding-level interpretability have been systematically analyzed, offering insights into the mechanisms driving decision-making processes.
@@
Previous research has examined the internal information distribution within these models Basu et al. (2024), highlighted the limitations arising from pre-trained vision encoders Tong et al. (2024), and investigated the issue of hallucinations in VLMs Bai et al. (2024); Jiang et al. (2025). Studies such as Huo et al. (2024) delve into neuron-level interpretations, uncovering how specific neurons contribute to multimodal reasoning. Similarly, Jiang et al. (2024) address hallucination issues in VLMs by examining the influence of vision-language representations. Both works highlight the importance of dissecting model components to improve transparency and robustness.
@@
On the methodological front, attention-based mechanisms remain a cornerstone for interpretability. Tao et al. (2024) explore global and local semantic representations, while Liu et al. (2023b) focus on visual instruction tuning to enhance alignment in cross-modal embeddings. Chefer et al. (2021) provides a technique to highlight the input pixels in an image that affects the decision of a transformer. Chen et al. (2022) further extends this approach for CLIP and Yellinek et al. (2023) suggests using a tree structure to improve the use of HilaCAM Chefer et al. (2021) in VLMs.
@@
Finally, causal approaches and model manipulation techniques have gained traction for evaluating VLMs. Palit et al. (2023) introduce causal tracing tools to understand information flow in vision-language systems, while Zhang et al. (2024) analyze information redundancy and relevance across reasoning tasks. These works demonstrate the efficacy of manipulating hidden states and attention mechanisms to investigate and improve model performance.
@@
Some datasets probe how well VLMs handle knowledge-intensive queries Lerner et al. (2024); Zhu et al. (2024); Onoe et al. (2025); Yanuka et al. (2025); Wei et al. (2025). For instance, InfoSeek Chen et al. (2023) is a VQA dataset tailored for information-seeking questions that cannot be resolved using only common sense knowledge. Their findings reveal that state-of-the-art VLMs often struggle with these questions; however, fine-tuning on InfoSeek enables them to leverage more fine-grained knowledge learned during pre-training. Unlike InfoSeek, which focuses on direct information-seeking queries, our dataset explicitly includes a bridge entity as part of a two-hop reasoning process—requiring both identification of the entity and extraction of the relevant fact. This design allows us to disentangle the efficacy of visual recognition from the subsequent factual retrieval and more directly analyze the gap between textual and visual representations.
@@
Our work is most related to Neo et al. (2025); Kaduri et al. (2024) who investigated processing of information in VLMs. Neo et al. (2025) provided insights into the role of attention layers in capturing visual attributes and object-specific details. Contemporary to us, Kaduri et al. (2024) demonstrated that query tokens store high-level image descriptors. Their findings reveal that middle layers of the model are critical for cross-modal knowledge transfer, while fine-grained visual details are extracted in a spatially localized manner. Yet, both of these works do not study the gap we reveal between the processing of textual and visual representations.
@@
Parcalabescu and Frank (2023) introduce a method that uses Shapley values to quantify how VLMs balance information from each modality, revealing varying degrees of unimodal collapse. Parcalabescu and Frank (2025) examine how VLM decoders rely on visual and textual inputs differently when producing answers versus explanations, noting greater reliance on text overall but an increased role for images in explanation tasks. Unlike these approaches, which focus on measuring how images and text are combined in general, our work studies a specific two-step reasoning process—entity identification followed by factual retrieval—and highlights a distinct performance gap when relying on visual instead of textual cues. Zhu et al. (2024) propose a pipeline to detect and mitigate what they refer to as parametric knowledge conflicts in large vision-language models, introducing a dynamic contrastive decoding method that improves performance on InfoSeek. While their work addresses cross-modal inconsistencies by focusing on parametric conflicts, we contribute a distinct perspective centered on bridging entity identification and factual retrieval.
@@
Interestingly, when compared to human behavior, the behavior of VLMs we find in this paper contrasts with findings from cognitive psychology. Specifically, humans tend to perform better with visual rather than textual information, a phenomenon known as the picture superiority effect (Jenkins et al., 1967). While explanations for this effect are still debated, the dual-coding theory (Paivio and Csapo, 1973) attributes this to images being encoded both visually and verbally, unlike text which is only encoded verbally. These findings suggest that humans benefit more from visual input, in contrast to VLMs, which are likely biased toward text due to its prevalence in their training data. We leave a further exploration of the relation between the two to future research.
@@ Appendix A Motivating Experiment - Results Breakdown By Type
Brands $1.07\times 10^{-21}$ $1.16\times 10^{-16}$ $4.78\times 10^{-21}$
@@ Appendix B Additional Experiment: Critical Information is Propagated Late
To investigate the stages of the model’s processing pipeline where entity identification occurs, we ask the model to identify the subject in an image and employ an attention knockout strategy following the methodology of Geva et al. (2023). This involves selectively prohibiting attention from textual tokens to image tokens at various layers of the model, effectively blocking cross-modal information flow. We do this by inserting $-inf$ in the attention mask at $[x,y]$ where $x$ are the indices of $\mathbf{h}_{1,t}^{\ell}\dots\mathbf{h}_{m,t}^{\ell},\mathbf{h}_{1,g}^{\ell}\dots\mathbf{h}_{k,g}^{\ell}$, $y$ are the indices of $\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell}$, and $\ell$ is any layer we want to knock out. We conduct this experiment in two directions:
@@
Notably, part of this experiment has been done in concurrent work by (Kaduri et al., 2024), but on different data, and asking different questions. This experiment checks whether the VLM trend follows Geva et al. (2023) which analyze the information flow for factual recall given textual inputs.
@@
These trends mostly agree with Geva et al. (2023) that knock out attention edges of subject tokens in LLMs. This may indicate that the model treats the image tokens similarly to subject textual tokens. One must remember that unlike the textual tokens, the image tokens are already processed and enriched within the vision encoder. This means that although one could hope that image tokens would enjoy extraction at much earlier stages, allowing for contextualization by the text tokens through more layers, the model exhibits suboptimal integration of image information, relying heavily on mid-layer processing while under utilizing earlier layers for cross-modal fusion.

### v3

@@ 1 Introduction
Over the past several years, vision-language models (VLMs) have made remarkable progress in linking visual inputs with natural language descriptions, enabling them to perform tasks such as image recognition clip; mpvr, video understanding lin2023video, and Visual Question Answering (VQA) liu2023visual; llava-next. Much of the existing research evaluating these models focuses on their capacity to understand and describe information presented in the input image: counting objects, assessing complex visual arrangements, and understanding interactions between humans and objects dang2024towards; Basu2024. While this has demonstrated that VLMs can integrate visual and textual information effectively, it does not fully address a critical question:
@@
Our empirical findings reveal a large accuracy drop when relying on visual representations rather than textual ones. Further probing suggests that this stems from the model investing too many layers in identifying the subject visually, leaving fewer layers for the second hop of factual reasoning (biran-etal-2024-hopping). To investigate these dynamics more concretely, we utilize techniques of mechanistic interpretability such as activation patching (NEURIPS2020_92650b2e) and attention knockout (geva-etal-2023-dissecting). They allow us to intervene in the model’s processing mid-inference, altering how and when image information influences the final answer. By tracking the flow of information through the network, we show that entity recognition and knowledge extraction are deeply entangled processes, with identification dominating the representational space longer than what one would expect. Consequently, the model’s internal knowledge is leveraged less efficiently.
@@ 2 Problem Setup
VLMs are architectures that integrate visual and textual modalities to perform tasks such as image captioning, visual question answering and more. Our VLM analysis in this work is mainly focused on the common architecture of LLaVA liu2023visual, and thus, we will follow the same notation.
@@
Every token in this input can be written as $\mathbf{h}_{i,v}^{0}$ or $\mathbf{h}_{i,t}^{0}$ where the $v$ and $t$ indicate a visual or textual token, the $i$ indicates the $i^{\text{th}}$ visual or textual tokens, and the 0 indicates input to layer 0. As the language model processes the entire input at each layer $l$, it builds a representation $\left(\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell},\mathbf{h}_{1,t}^{\ell}\dots\mathbf{h}_{m,t}^{\ell},\mathbf{h}_{1,g}^{\ell}\dots\mathbf{h}_{k,g}^{\ell}\right)$, where $\mathbf{h}_{1,g}^{\ell}...\mathbf{h}_{k,g}^{\ell}$ are the generated tokens hidden representations, $\left|\mathbf{H}_{\mathbf{v}}\right|=n$, $\left|\mathbf{H}_{\mathbf{t}}\right|=m$ and $\ell$ is the layer. We refer to the tokens once inside the LLM as hidden states.
@@
Through activation patching (NEURIPS2020_92650b2e) and other mechanistic interpretability methods, we trace how and where entity recognition occurs, and how it impacts the model’s ability to retrieve the correct facts. Section 5 describes how we employ activation patching to identify the layer at which the model extracts the entity identity, establishing an upper bound on identification layers. Building on these findings, Section 6 applies forward activation patching to investigate whether the entity identity could be inferred at earlier layers.
We include additional experiments using attention knockouts (geva-etal-2023-dissecting) in appendix B, which further validate our conclusions. Throughout the paper, we use POPVQA to systematically evaluate VLM performance on visually driven factual retrieval and demonstrate how visual and textual components interact within a unified transformer architecture.
@@ 3 The POPVQA Dataset
In order to isolate the model’s ability to extract knowledge about a given entity from its ability to identify it in the first place, we want to restrict our questions to entities that the model recognizes. Other datasets that have knowledge seeking goals such as infoSeek (chen-etal-2023-pre-trained) or VeQuAE (lerner2022viquae) do not include the entity in the image but rather only the answer to the question, thus, making it difficult to do this separation. Therefore, we built the POPVQA dataset to facilitate this need. It is inspired by PopQA mallen-etal-2023-trust that created a QA dataset for various entities using Wikipedia.
The first part of POPVQA comprises of 15,395 entity-image pairs from Wikidata (vrandevcic2014wikidata) along with aliases and popularity score (number of page views). The vast majority of entities are celebrities (63.6%). To allow ablation of other entity types, we added landmarks (18%), such as famous structures or UNESCO sites, famous paintings (14.6%), and brand logos (3.8%). All images are resized to 336 × 336 which is a standard input resolution, by adding black borders to create a square, then resizing. This was done in order to maintain the original aspect ratio, without losing any details. The popularity score was used only in the creation of the dataset such that we select only entities with high popularity, which increases the chance that the models will recognize them mallen-etal-2023-trust. We do not use this value directly in the experiments.
The second part of POPVQA provides information seeking questions about the entities; at least two per entity. To generate the questions, we convert subject-relation-object triplets from Wikidata into question-answer pairs using manually designed templates (e.g., petroni-etal-2019-language; sciavolino-etal-2021-simple). For example, the triplet $(\textit{Robin Williams, spouse, Marsha Garces})$ is transformed into the question, “Who is the spouse of Robin Williams?”, with the answer “Marsha Garces”. We filter out questions where the answer can be deduced from the image alone without recognizing the entity, such as eye color. We point out that this is not bulletproof as, for example, the native language of an athlete can be guessed if the athlete is wearing a national team jersey.
@@ 4 Motivating Experiment
We show results for the following models; LLaVA-1.5-7B (liu2023visual) as the canonic base VLM; we refer to it as LLaVA7B. LLaVA-MORE-8B (cocchi2025llavamorecomparativestudyllms) which uses a larger Llama 3.1 as its LLM and we try both variants using CLIP (Radford2021LearningTV) and SigLIP (10377550) as backbones allowing for controlled comparison between visual backbones; we refer to these as LM-CLIP and LM-SigLIP. We use LLaVA-1.6-34B to assess the effect of model scale, and refer to it as LLaVA34B. We also compare to Qwen2-VL (Qwen2-VL) that has a different architecture and training scheme.
@@ 7 Related Work
While VLMs demonstrate impressive capabilities in tasks such as visual reasoning, object recognition, and multimodal understanding Basu2024, their internal mechanisms and interpretability remain active areas of research. Below, we review the related works categorized into general VLM interpretability, cross-modal information flow, attention mechanisms, and mechanistic approaches to visual token processing.
@@
dang2024towards survey VLMs proposing a framework categorizing interpretability across data, model, and training perspectives. They highlight the alignment and decomposition of modalities as key challenges, emphasizing the need for explainability to ensure trustworthiness in multimodal AI applications. In this context, techniques for token-level, neuron-level, and embedding-level interpretability have been systematically analyzed, offering insights into the mechanisms driving decision-making processes.
@@
Previous research has examined the internal information distribution within these models Basu2024, highlighted the limitations arising from pre-trained vision encoders Tong2024, and investigated the issue of hallucinations in VLMs Bai2024; jiang2025interpreting. Studies such as huo2024mmneuron delve into neuron-level interpretations, uncovering how specific neurons contribute to multimodal reasoning. Similarly, jiang2024interpreting address hallucination issues in VLMs by examining the influence of vision-language representations. Both works highlight the importance of dissecting model components to improve transparency and robustness.
@@
On the methodological front, attention-based mechanisms remain a cornerstone for interpretability. tao2024probing explore global and local semantic representations, while liu2023visual focus on visual instruction tuning to enhance alignment in cross-modal embeddings. Chefer_2021_ICCV provides a technique to highlight the input pixels in an image that affects the decision of a transformer. Chen_2022_ACCV further extends this approach for CLIP and yellinek20233vlusingtreesteach suggests using a tree structure to improve the use of HilaCAM Chefer_2021_ICCV in VLMs.
@@
Finally, causal approaches and model manipulation techniques have gained traction for evaluating VLMs. palit2023towards introduce causal tracing tools to understand information flow in vision-language systems, while zhang2024from analyze information redundancy and relevance across reasoning tasks. These works demonstrate the efficacy of manipulating hidden states and attention mechanisms to investigate and improve model performance.
@@
Some datasets probe how well VLMs handle knowledge-intensive queries Lerner2024Cross-Modal; zhu2024unravelingcrossmodalityknowledgeconflicts; Onoe2025DOCCI; yanuka-etal-2025-bridging; Wei2025UniIR. For instance, InfoSeek chen-etal-2023-pre-trained is a VQA dataset tailored for information-seeking questions that cannot be resolved using only common sense knowledge. Their findings reveal that state-of-the-art VLMs often struggle with these questions; however, fine-tuning on InfoSeek enables them to leverage more fine-grained knowledge learned during pre-training. Unlike InfoSeek, which focuses on direct information-seeking queries, our dataset explicitly includes a bridge entity as part of a two-hop reasoning process—requiring both identification of the entity and extraction of the relevant fact. This design allows us to disentangle the efficacy of visual recognition from the subsequent factual retrieval and more directly analyze the gap between textual and visual representations.
@@
Our work is most related to neo2025towards; kaduri2024whats who investigated processing of information in VLMs. neo2025towards provided insights into the role of attention layers in capturing visual attributes and object-specific details. Contemporary to us, kaduri2024whats demonstrated that query tokens store high-level image descriptors. Their findings reveal that middle layers of the model are critical for cross-modal knowledge transfer, while fine-grained visual details are extracted in a spatially localized manner. Yet, both of these works do not study the gap we reveal between the processing of textual and visual representations.
@@
parcalabescu-frank-2023-mm introduce a method that uses Shapley values to quantify how VLMs balance information from each modality, revealing varying degrees of unimodal collapse. parcalabescu2025do examine how VLM decoders rely on visual and textual inputs differently when producing answers versus explanations, noting greater reliance on text overall but an increased role for images in explanation tasks. Unlike these approaches, which focus on measuring how images and text are combined in general, our work studies a specific two-step reasoning process—entity identification followed by factual retrieval—and highlights a distinct performance gap when relying on visual instead of textual cues. zhu2024unravelingcrossmodalityknowledgeconflicts propose a pipeline to detect and mitigate what they refer to as parametric knowledge conflicts in large vision-language models, introducing a dynamic contrastive decoding method that improves performance on InfoSeek. While their work addresses cross-modal inconsistencies by focusing on parametric conflicts, we contribute a distinct perspective centered on bridging entity identification and factual retrieval.
@@
Interestingly, when compared to human behavior, the behavior of VLMs we find in this paper contrasts with findings from cognitive psychology. Specifically, humans tend to perform better with visual rather than textual information, a phenomenon known as the picture superiority effect (jenkins1967differential). While explanations for this effect are still debated, the dual-coding theory (PAIVIO) attributes this to images being encoded both visually and verbally, unlike text which is only encoded verbally. These findings suggest that humans benefit more from visual input, in contrast to VLMs, which are likely biased toward text due to its prevalence in their training data. We leave a further exploration of the relation between the two to future research.
@@ Appendix A Motivating Experiment - Results Breakdown By Type
Brands $07\times 10^{-21} 1.1$$16\times 10^{-16}$ $4.78\times 10^{-21}$
@@ Appendix B Additional Experiment: Critical Information is Propagated Late
To investigate the stages of the model’s processing pipeline where entity identification occurs, we ask the model to identify the subject in an image and employ an attention knockout strategy following the methodology of geva-etal-2023-dissecting. This involves selectively prohibiting attention from textual tokens to image tokens at various layers of the model, effectively blocking cross-modal information flow. We do this by inserting inf-inf in the attention mask at $[x,y]$ where $x$ are the indices of $\mathbf{h}_{1,t}^{\ell}\dots\mathbf{h}_{m,t}^{\ell},\mathbf{h}_{1,g}^{\ell}\dots\mathbf{h}_{k,g}^{\ell}$, $y$ are the indices of $\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell}$, and $\ell$ is any layer we want to knock out. We conduct this experiment in two directions:
@@
Notably, part of this experiment has been done in concurrent work by (kaduri2024whats), but on different data, and asking different questions. This experiment checks whether the VLM trend follows geva-etal-2023-dissecting which analyze the information flow for factual recall given textual inputs.
@@
These trends mostly agree with geva-etal-2023-dissecting that knock out attention edges of subject tokens in LLMs. This may indicate that the model treats the image tokens similarly to subject textual tokens. One must remember that unlike the textual tokens, the image tokens are already processed and enriched within the vision encoder. This means that although one could hope that image tokens would enjoy extraction at much earlier stages, allowing for contextualization by the text tokens through more layers, the model exhibits suboptimal integration of image information, relying heavily on mid-layer processing while under utilizing earlier layers for cross-modal fusion.